WA EUP In manufacturing its products, Trevano Corp. adds all direct material at the beginning of...

WA EUP
In manufacturing its products, Trevano Corp. adds all direct material at the beginning of the production process. The company’s direct labor and overhead are considered to be continuously at the same degree of completion. September production information is as follows:

In manufacturing its products, Trevano Corp. adds all direct material at the beginning of the production process. The company’s direct labor and overhead are considered to be continuously at the same degree of completion. September production information is as follows:

Beginning WIP Inventory 10,000 pounds
Started during September 25,000 pounds
Completed during September 32,000 pounds

As of September 1, the beginning WIP Inventory was 30 percent complete as to labor and overhead. On September 30, the ending WIP Inventory was 80 percent complete as to conversion.

a. Determine the total number of pounds to account for if Trevano uses the weighted average process costing method.
Answer pounds
b. Determine the equivalent units of production for direct material. Answer pounds
c. Determine the equivalent units of production for direct labor and overhead. Answer pounds
Answer #1

image.png

answered by: qbank
Answer #2

Total number of pounds to account for= 35000 units.

A Beginning WIP                   10,000
B Introduced                   25,000
C=A+B TOTAL units accounted for                   35,000

.

b. Determine the equivalent units of production for direct material. =35,000 pounds
c. Determine the equivalent units of production for direct labor and overhead. = 34,400 pounds

Working

Reconciliation of Units
A Beginning WIP                   10,000
B Introduced                   25,000
C=A+B TOTAL                   35,000
D Transferred out                   32,000
E=C-D Ending WIP                     3,000

.

Statement of Equivalent Units(Weighted average)
Material Labor and overhead
Units Complete % Equivalent units Complete % Equivalent units
Transferred out              32,000 100%                    32,000 100%                   32,000
Ending WIP                3,000 100%                       3,000 80%                     2,400
Total              35,000 Total                    35,000 Total                   34,400